OpenStudy (anonymous):

a=-2 , b=-1 , c=4 , d=3 Solve) √2a-3b+4=?

OpenStudy (callisto):

\[\sqrt{2a-3b+4}\] Is this your question?

OpenStudy (anonymous):

put the values of a and b..then solve

OpenStudy (anonymous):

no its 2a(then the square root of 2) +b-3c

OpenStudy (callisto):

\[\sqrt {2}a +b - 3c\] ?

OpenStudy (anonymous):

yah

OpenStudy (callisto):

Put a = -2 b=-1 and c=4 into that. \[\sqrt2a + b -3c = \sqrt 2 (-2) + (-1) -3(4) =?\]